OpenStudy (sleepyjess):

@ganeshie8 one more please?

OpenStudy (sleepyjess):

\(\sf\dfrac{sinx}{1-cosx}+\dfrac{sinx}{1+cosx}=2cscx\)

OpenStudy (anonymous):

can I help??

OpenStudy (turingtest):

simplify the left side by combining the fractions

OpenStudy (anonymous):

take sinx common

OpenStudy (sleepyjess):

\(\sf\dfrac{2sinx}{2}=2cscx\)?

OpenStudy (anonymous):

solve \[1/1-\cos x+1/1+cosx\]

OpenStudy (turingtest):

how did you get a 2 on the bottom?

OpenStudy (anonymous):

sinx comes in denominator

OpenStudy (anonymous):

small mistake @sleepyjess

Parth (parthkohli):

Welp, looks like you've asked the same question again.

OpenStudy (anonymous):

on solving u will get 2/sin^2x

OpenStudy (sleepyjess):

1+1 in the denominators

OpenStudy (turingtest):

(1-cosx)(1+cosx)=1-cos^2x

OpenStudy (anonymous):

\[finally~u~have~ 2sinx/\sin^2x\]

OpenStudy (anonymous):

so sinx comes in the denominator.......

OpenStudy (anonymous):

which gives us the RHS

OpenStudy (anonymous):

1+1 in the Nr